yourgrandpa:

members of a band estimated that 200 people were at their show at night. their manager told them that 227 people had actually been at the show. what was the band's percent error?

AZ:

\( \text{% error} = \dfrac{|\text{actual - expected}|}{\text{actual}} \times 100\)

AZ:

How many people actually showed up and how many were estimated/expected to be there?

yourgrandpa:

227 showed up and 200 expected to be there ???

AZ:

yes, so let's put those numbers into the equation \( \text{% error} = \dfrac{|\text{227- 200}|}{\text{227}} \times 100\) 227 - 200 = ?

yourgrandpa:

27 ??

AZ:

Good | | is just absolute value and it's in case we get a number number to make it positive Now what is 27/227 = ?

yourgrandpa:

227/27 ??

NottheMonga2087:

divided

NottheMonga2087:

227/27=

yourgrandpa:

oo ok

yourgrandpa:

0.118943

AZ:

\( \text{% error} = \dfrac{|\text{227- 200}|}{\text{227}} \times 100\) \( \text{% error} = \dfrac{27}{\text{227}} \times 100\) We have to simplify the fraction so divide it

AZ:

That's good! Now multiply it by 100 and that's your % error

yourgrandpa:

11.894300

AZ:

There you go!

yourgrandpa:

so whats the answer ???? lol

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa 11.894300 \(\color{#0cbb34}{\text{End of Quote}}\) You got it

AZ:

Your question doesn't specify how much they want you to round by

AZ:

So maybe 11.89% is fine

NottheMonga2087:

Good Job ;]

yourgrandpa:

oo thank you so much(: